(Y2+z2-x2)p-2xyq+2xz=0


1
Expert's answer
2021-05-07T09:40:39-0400

Given equation is Lagrange's linear equation

Pp+Qq=R

where P=y2+z2-x2, Q=-2xy, R=-2xz.

The auxiliary equation is

"\frac{dx}{P}=\frac{y}{Q}=\frac{dz}{R}"

"\frac{dx}{y^2+z^2-x^2}=\frac{y}{-2xy}=\frac{dz}{-2xz}"....................................................(1)

Taking last two ratios,

"\frac{dy}{-2xy}=\frac{dz}{-2xz}\newline
\frac{dy}{y}=\frac{dz}{z}\newline

\text{Integrating both side,}\newline
logy=logz+logc,\space \text {where c is a constant.}\newline
c=\frac{y}{z}"


Taking Lagrangian multipliers as, x,y,z, each ratios of (1),

"\frac{xdx+ydy+zdz}{-x(x^2+y^2+z^2)}"


Now take,

"\frac{dy}{-2xy}=\frac{xdx+ydy+zdz}{-x(x^2+y^2+z^2)}\newline

\frac{dy}{y}=\frac{d(x^2+y^2+z^2)}{(x^2+y^2+z^2)}\newline

\text{Integrating both side,}\newline

logy=log(x^2+y^2+z^2)+logb\newline

∴\frac{y}{x^2+y^2+z^2}=b\newline

∴ 
\text{The general equation is}\newline

ϕ(\frac{y}{z},\frac{y}{x^2+y^2+z^2})=0"